brandi1986:

Evaluate 9 P 4. 15,120 36 3,024 504

brandi1986:

can someone help me with the answer to this

brandi1986:

hello

Logic007:

\(^nP_r=\frac{(n!)}{(n-r)!}\)

Logic007:

so \(^9P_4=\frac{(9!)}{(9-4)!}\)

Logic007:

\(^9P_4=\frac{9!}{5!}=\frac{9\times8\times7\times6\times5!}{5!}\)

Logic007:

Then,simplify it \(=\frac{9\times8\times7\times6\times\cancel{5!}}{\cancel{5!}}\)

Logic007:

=? solve it! @brandi1986